# Hai-Long Sun1,2,3   Da-Wei Zhou1,2  Yang Li3  Shiyin Lu3  Chao Yi1,2 Qing-Guo Chen3  Zhao Xu3  Weihua Luo3  Kaifu Zhang3  De-Chuan Zhan1,2  Han-Jia Ye1,2

### v2

@@ 4.2 Pilot Study
To address the challenge of multilingual erosion in MLLMs due to the dominance of English in image-text data, we hypothesize that there is an inherent mismatch between visual tokens ${\bf H}_{v}$ and textual tokens ${\bf H}_{t}$, which tends to bias them towards English semantics, making the model more likely to generate outputs in English. Specifically, the widely-used vision encoder of OpenAI-CLIP [51] is pre-trained on a vast corpus of English-centric image-text pairs, resulting in visual representations more aligned with English.
To explore this phenomenon, we train two distinct models using the same data: one incorporating OpenAI-CLIP vision encoder and the other integrating Chinese-CLIP [62] vision encoder. As shown in Figure 1, the model equipped with OpenAI-CLIP struggles to generate suitable outputs according to the Chinese inputs, whereas the model using Chinese-CLIP not only understands the queries but also generates appropriate outputs in Chinese. Moreover, to further evaluate the model’s Chinese capability, we test it on Chinese datasets and observe improved performance: from 66.4 to 68.3 on MMBench-CN and from 62.4 to 66.1 on MMMB-zh when utilizing Chinese-CLIP.
@@ 4.3 Textual Guidance to Drive Visual Token Alignment
Due to the low-resource nature of non-English multimodal data (e.g., lack of large-scale, high-quality image-text data), we need to use nearly the same amount of data as LLaVA to enhance the model’s multilingual capabilities. Furthermore, motivated by these interesting findings and aiming to align visual tokens with textual embeddings at the language level, we propose Parrot, a novel approach that leverages textual guidance to facilitate the multilingual alignment of visual features. Parrot enables the transition of English-biased visual features acquired through the OpenAI-CLIP to accommodate other languages. This approach ensures that it can provide language-specific visual tokens to LLM based on the multiple language inputs, thereby enhancing its multilingual capabilities.
@@
First, we extract visual features through the vision encoder and transform them into language embedding tokens ${\bf H}_{v}$ using a projector. We obtain the embeddings ${\bf H}_{t}\in\mathbb{R}^{N\times C}$ derived from text inputs via the word embedding table. Subsequently, to convert the English-biased features into language-specific features using textual guidance, we employ a cross-modal cross-attention mechanism to obtain ${\bf H}_{v}^{\prime}\in\mathbb{R}^{C}$:
@@
where $\mathbf{Q}$ equals the matrix ${\bf H}_{v}$, $\mathbf{K}$ and $\mathbf{V}$ are equivalent to ${\bf H}_{t}$. ${\bf H}_{v}^{\text{cls}}\in\mathbb{R}^{C}$ is the [CLS] token of ${\bf H}_{v}$. This process allows the visual features to be dynamically adjusted based on the multilingual inputs.
Since the projected language embedding tokens ${\bf H}_{v}$ are English-biased, we need to convert them to language-specific embeddings for different languages. To this end, we introduce a lightweight Mix-of-Experts (MoE) module, which includes a router and several language transformation experts. The router of MoE is a linear layer that generates a probability distribution over the set of experts $\mathcal{E}=[e_{1},e_{2},\cdots,e_{E}]$, effectively predicting the probability of selecting and activating each expert. Each expert is an MLP designed to convert English-biased embeddings into language-specific embeddings. The inputs to experts $\mathcal{E}$ is ${\bf H}_{v}$, and the outputs have the same dimensions as the inputs.
@@
which selects and activates the specific experts. Moreover, we process the English-biased embeddings ${\bf H}_{v}$ through the selected experts to convert them into language-specific visual representations:
@@
This approach effectively aligns English-biased embeddings with multiple languages, ensuring a more accurate and comprehensive representation across different linguistic contexts. To stabilize training and convert visual embeddings with less variance in visual-semantic information, ensuring the model performs well in tasks beyond the multilingual multimodal domain, we utilize MoE reweighting to obtain the final language-specific visual embeddings ${\bf G}_{v}$:
@@
where $\alpha$ is the trade-off parameter. In conclusion, we first fuse the visual and textual inputs via Eq. 2 to transform the visual embeddings with textual guidance. Moreover, the fused result is inputted into the MoE module to select and activate the most relevant language experts via Eq. 3 and then obtain the language-specific embeddings via Eq.  4. Lastly, we employ MoE reweighting to convert visual embeddings with less variance in original visual-semantic information 5. This approach enables us to endow the MLLM with multilingual capabilities using as little multilingual data as possible. Figure 4 illustrates the architecture, the detailed MoE module, and the training stages of Parrot.
Table 1: Details on the Parrot’s training data, derived from publicly available datasets and our in-house multilingual data.
Training Stage Datasets Samples Total
Stage 1 LLaVA-1.5-pretrain [38] 558K 1.2M
Laion-Caption [55] 12K
CC12M-Caption [9] 645K
Stage 2 LLaVA-1.5-finetune [38] 665K 793K
ShareGPT4V-zh [12] 71K
ShareGPT4V-pt [12] 14K
ShareGPT4V-ar [12] 12K
ShareGPT4V-tr [12] 17K
ShareGPT4V-ru [12] 14K
@@ 4.4 Training Stage
Our goal is to utilize as little multilingual data as possible to enhance the multilingual capabilities of MLLMs. The whole training procedure is divided into two distinct stages:
Stage 1: Modality Alignment. In this stage, we keep both the vision encoder and the LLM weights frozen, focusing solely on optimizing the projectors to align the visual features ${\bf H}_{v}$ with the pre-trained LLM word embedding. This stage can be likened to training a visual tokenizer that is compatible with the frozen LLM. To enhance the diversity of images, we extract a portion of data from LAION [55] and CC12M [9] datasets and construct the in-house caption data through GPT-4V.
Stage 2: Instruction Tuning for Multilingual Alignment. We still keep the vision encoder weights frozen while continuing to train the projector, MoE, and LLM. Due to the design of the MoE module, Parrot can rapidly learn to align visual representations across multiple languages by using a small amount of multilingual image-text data. As shown in Table 1, we only use nearly 10K training data for each language in stage 2. This approach is particularly beneficial given the scarcity of data resources in low-resource languages.
To address the challenge of limited data in non-English languages, we use a semi-automatic approach to acquire image-text data. Initially, we partition the ShareGPT4V dataset [12] randomly for each language, extracting a selection of non-duplicate, non-parallel image-text data for training. Subsequently, we implement a translation and calibration scheme using GPT-4 to convert English texts into texts of other languages. Recognizing that this step may introduce noise and potential translation errors, we apply a manual calibration process to further fine-tune and clean the data, thereby obtaining high-quality multilingual image-text data.
This two-stage training approach ensures effective modality and multilingual alignment, even with limited Non-English data, aligning well with the realities of data scarcity in low-resource languages.
@@ 5 Experiments
In this section, we begin with an overview of the experimental framework, providing details on specific implementations, evaluation benchmarks, and MLLMs used for comparative evaluation. Following this, we conduct a comprehensive comparison of Parrot with the state-of-the-art approaches using multilingual benchmarks. Additionally, we compare Parrot with leading models across a range of multimodal tasks. Finally, this section concludes with ablation studies and visualization of multilingual cases, highlighting the exceptional ability of Parrot in handling multilingual tasks.
5.1 Experimental Setup
Implementation Details: In this study, we configure Parrot with the pre-trained CLIP ViT-L/14 [51] as the vision encoder and Qwen1.5-Chat [5] as the backbone for LLM. The initial learning rates for the two stages are set at $1e^{-3}$ and $2e^{-5}$, respectively, with the batch size of 256 and 128. The entire training process is notably optimized to 21 hours on the 16$\times$×A100 GPUs setup, attributed to the use of the relatively small training datasets. Additionally, BF16 and TF32 precision formats are employed to meticulously balance speed and accuracy throughout the training process. As defined in Eq. 4, we set the number of experts to six to match the number of languages. Each expert is an MLP composed of two linear layers with SiLU [19] activation function. More details are shown in Table 5.
Evaluation Benchmark: Our evaluation is divided into two parts: one evaluates the multilingual capabilities of MLLMs, while the other assesses its overall performance. The first evaluation is performed on two datasets: MMBench [39] and a newly developed benchmark MMMB. For MMBench, we expand it to include six languages through translation via GPT-4, followed by manual verification. For MMMB, we construct it following the methodology described in Section 3. We present the accuracy for each language in Table 2. Furthermore, the second evaluation covers a wide broad range of multimodal tasks, such as MME [22], MMStar [11], ScienceQA [41], RealWorldQA [61] and SEED-Bench [28], with performance reported using a radar chart in Figure 5(b).
Comparison Models: For comprehensive comparisons, we select leading open-source models in MLLMs, including LLaVA-1.5 [29], LLaVA-NeXT [37], Qwen-VL [6], Monkey [35], VisualGLM [18], VisCPM [24], MiniGPT-4-v2 [75], shareGPT4V [12], InstructBLIP [17], mPLUG-Owl2 [64], Mini-Gemini [34]. Furthermore, we incorporate closed-source methods in our benchmarks, including GPT-4V [12], Qwen-VL-MAX [6], and Gemini Pro [54], to demonstrate their remarkable performance. For the evaluation process, we employ the VLMEvalKit in OpenCompass [16], ensuring consistent configuration settings across all methods to maintain fairness in comparison. For most of the mentioned methods, we directly use the VLMEvalKit implementation. Alternatively, we integrate other methods not officially provided into this framework to ensure consistency in evaluation.
5.2 Main Results
In this section, we present the results of the multilingual experiment in Table 2 and the overall experiment in Figure 5(b). According to Table 2, Parrot-14B achieves state-of-the-art (SOTA) performance in all languages on the MMBench benchmark and also achieves the SOTA performance in five languages on the MMMB benchmark, with English in the second place. The multilingual capabilities of Parrot-14B closely reach that of GPT-4V, demonstrating the exceptional ability of our proposed architecture. Notably, Parrot-7B also validates the SOTA performance on both benchmarks across all languages, even surpassing the LLaVA-NeXT-13B. Additionally, as shown in Figure 5(b), this evaluation aims to show that Parrot not only possesses excellent multilingual capabilities but also provides an overall understanding of Parrot’s capabilities in handling various complex multimodal tasks (e.g., MME [22], MMStar [11], and SEED-Bench [28]). Additionally, as depicted in Figure 5(c), we visualize the expert distributions within the MoE. It is evident that the second expert is predominantly activated when using the Chinese prompt, indicating that different experts are utilized for various language prompts. In existing multilingual benchmarks, Parrot also demonstrates competitive performance while using less than 1% of the data compared to other multilingual MLLMs, as illustrated in Figure 6.
Table 2: Accuracy performance comparison on multilingual benchmarks. We report all compared methods with VLMEvalKit [16]. The best and second results are shown in bold and underline, respectively.
@@ 5.2 Main Results
Open-source models
LLaVA-1.5 [36] Vicuna-v1.5-7B 67.07 58.83 59.76 43.50 46.43 59.06 65.37 58.33 59.02 36.16 43.90 56.95
LLaVA-1.5 [36] Vicuna-v1.5-13B 69.76 62.86 60.76 45.49 54.44 62.69 68.98 63.23 62.97 46.56 53.17 61.59
LLaVA-NeXT [37] Vicuna-v1.5-7B 70.87 61.57 61.81 42.74 46.95 63.85 67.95 60.56 60.39 38.40 45.36 59.62
LLaVA-NeXT [37] Vicuna-v1.5-13B 74.44 67.19 63.21 45.36 53.09 68.24 70.87 64.51 64.08 45.36 52.92 61.85
Qwen-VL [6] Qwen-7B 52.63 36.37 38.65 36.54 37.42 40.70 42.26 22.25 25.08 18.72 26.37 28.17
Qwen-VL-Chat [6] Qwen-7B 56.02 57.77 46.37 43.04 41.05 48.65 54.29 56.52 43.12 35.73 39.17 42.86
MiniGPT-4-v2 [75] LLaMA2-13B 38.71 30.05 31.52 26.60 26.02 29.23 23.88 11.76 14.26 2.49 6.78 12.54
ShareGPT4V [12] Vicuna-v1.5-7B 69.24 60.23 60.29 43.57 45.26 61.23 69.59 61.6 59.62 37.37 43.38 59.45
InstructBLIP [17] Vicuna-7B 39.47 32.92 35.67 23.80 28.36 36.37 27.83 18.81 27.14 3.26 8.50 20.87
mPLUG-Owl2 [64] LLaMA2-7B 67.25 60.99 59.70 45.78 45.43 62.63 66.15 59.36 58.24 37.88 47.68 60.39
Monkey [35] Qwen-VL-7B 66.02 58.18 46.31 38.83 37.66 48.59 58.07 53.52 49.57 31.01 31.35 45.18
Monkey-chat [35] Qwen-VL-7B 71.63 66.54 60.35 48.77 46.31 58.59 70.79 65.72 65.03 46.90 48.10 59.36
VisualGLM [18] ChatGLM-6B 31.05 18.07 19.42 15.38 22.81 19.77 23.2 17.18 11.43 2.92 6.62 5.33
VisCPM-Chat [24] CPM-Bee-10B 53.10 47.54 28.19 26.90 26.78 26.84 45.88 46.39 15.81 1.46 9.19 1.20
@@
Parrot Qwen1.5-14B 73.92 71.64 69.82 68.13 64.33 70.18 74.40 72.25 69.16 66.15 64.52 69.33
Closed-source models
GPT-4V [46] Private 74.97 74.21 71.46 73.51 68.95 73.10 77.60 74.40 72.51 72.34 70.53 74.83
Gemini Pro [58] Private 75.03 71.87 70.64 69.94 69.59 72.69 73.63 72.08 70.27 61.08 69.76 70.45
Qwen-VL-MAX [6] Private 77.19 75.26 72.16 70.82 66.02 74.21 76.80 77.58 74.57 75.00 69.07 75.00
5.3 Ablation Study
Ablation study on each component. We conduct an ablation experiment on the multilingual data and the MoE module. As shown in Figure 5(a), using multilingual data improves performance in each language. Moreover, the MoE module significantly improves performance, demonstrating the effectiveness of our proposed method.

### v3

@@ 3.2 Pilot Study
To address the challenge of multilingual erosion in MLLMs due to the dominance of English in image-text data, we empirically observe an inherent mismatch between visual tokens ${\bf H}_{v}$ and textual tokens ${\bf H}_{t}$, which biases the model towards English semantics and increases the likelihood of English outputs. Specifically, the widely-used OpenAI-CLIP vision encoder (Radford et al., 2021), pre-trained on a large English-centric image-text corpus, yields visual representations more aligned with English.
To investigate this, we conduct the ablation study using OpenAI-CLIP and Chinese-CLIP (Yang et al., 2022). As shown in Figure 1, the OpenAI-CLIP model struggles with Chinese inputs, while the Chinese-CLIP model effectively processes and generates Chinese outputs. Additionally, to explore the distance between the visual features generated by different encoders and the Chinese prompts, we plot t-SNE visualizations for a more intuitive understanding. As illustrated in Figure 6, the t-SNE further reveal that the visual features from Chinese-CLIP are more closely aligned with the Chinese prompts.
@@ 3.3 Textual Guidance to Drive Visual Token Alignment
First of all, we have to address a key challenge: due to the limited availability of non-English multimodal data, we cannot rely on extensive multilingual datasets to enhance the multilingual performance of MLLMs. Therefore, it is necessary to design a method that efficiently aligns visual and textual features at the language level, rather than relying on large multilingual data. Motivated by the pilot study, we aim to directly align visual tokens with textual embeddings at the language level. To this end, we propose Parrot, a novel method that leverages textual guidance to facilitate the multilingual alignment of visual features. Parrot enables the transition of English-biased visual features acquired through the OpenAI-CLIP to accommodate other languages, ensuring language-specific visual tokens are generated for LLMs based on multilingual inputs, thereby enhancing multilingual abilities.
@@
First, we extract visual features through the vision encoder and transform them into language embedding tokens ${\bf H}_{v}$ using a projector. We obtain the embeddings ${\bf H}_{t}\in\mathbb{R}^{N\times C}$ derived from text inputs via the word embedding table. Subsequently, to convert the English-biased features into language-specific ones, we employ a cross-modal cross-attention mechanism to obtain ${\bf H}_{v}^{\prime}\in\mathbb{R}^{C}$:
@@
where $\mathbf{Q}={\bf H}_{v}$, and both $\mathbf{K}$ and $\mathbf{V}$ are equivalent to ${\bf H}_{t}$. ${\bf H}_{v}^{\text{cls}}\in\mathbb{R}^{C}$ represents the [CLS] token of ${\bf H}_{v}$. This process allows the visual features to be dynamically adjusted and transformed into language-specific semantic embeddings based on multilingual inputs.
Since the projected language embedding tokens ${\bf H}_{v}$ are biased towards English, we need to convert them into language-specific embeddings for different languages. To this end, we introduce a lightweight Mixture-of-Experts (MoE) module, which consists of a router and several language transformation experts. The MoE router is a linear layer that generates a probability distribution over the set of experts $\mathcal{E}=[e_{1},e_{2},\cdots,e_{E}]$, effectively selecting and activating specific experts. Each expert is an MLP designed to convert English-biased embeddings into language-specific embeddings. The inputs to experts $\mathcal{E}$ is ${\bf H}_{v}$, and the outputs have the same dimensions as the inputs.
@@
which selects and activates the appropriate experts. Next, we process the English-biased embeddings ${\bf H}_{v}$ through the selected experts to convert them into language-specific visual representations:
@@
This approach effectively aligns English-biased embeddings with multiple languages, ensuring a more accurate and comprehensive representation across different linguistic contexts. To stabilize training and reduce the variance in visual-semantic information, ensuring the model performs well beyond the multilingual multimodal domain, we employ MoE reweighting to obtain the final language-specific visual embeddings ${\bf G}_{v}$:
@@
where $\alpha$ is a trade-off parameter. In conclusion, we first fuse the visual and textual inputs via Eq. 2 to transform the visual embeddings with textual guidance. The fused result is then input into the MoE module, which selects and activates the most relevant language experts via Eq. 3 to obtain language-specific embeddings as shown in Eq. 4. Finally, MoE reweighting is applied to convert visual embeddings with less variance in original visual-semantic information 5. This approach enables the MLLM to gain multilingual capabilities using minimal multilingual data. Figure 5 illustrates the architecture, the detailed MoE module, and the training stages of Parrot.
@@ 3.4 Training Stage
Our goal is to enhance the multilingual capabilities of MLLMs with minimal multilingual data. The training procedure is divided into two distinct stages:
Stage 1: Modality Alignment. In this stage, we freeze both the vision encoder and the LLM weights, focusing solely on optimizing the projectors to bridge the modality gap. Notably, the MoE module is bypassed entirely, meaning the image tokens do not pass through the MoE since the primary goal of this stage is to train the projector using a large number of image-text pairs. This enables the projector to align image tokens and textual tokens effectively without interference from the untrained MoE module.
Stage 2: Instruction Tuning for Multilingual Alignment. We continue to freeze the vision encoder weights while training all other modules. In this stage, we introduce multilingual training data and randomly initialize the parameters of MoE. The MoE is optimized with textual guidance, which drives the alignment of visual tokens while leveraging the well-trained projector. The prior alignment achieved in the pre-training stage facilitates efficient optimization of the MoE during this phase. The entire training process of Parrot is outlined in pseudocode, as shown in Algorithm 1 in the Appendix.
To address data scarcity in non-English languages, we use a semi-automatic method similar to the one depicted in Figure 3 to acquire image-text data. We randomly partition the ShareGPT4V dataset (Chen et al., 2023b) for each language, extracting non-duplicate, non-parallel image-text pairs for training, ultimately obtaining nearly 10K samples per language. This two-stage training approach ensures effective modality and multilingual alignment, even with limited non-English data, aligning well with the challenges of data scarcity in low-resource languages.
@@ 4 Experiments
In this section, we begin with an overview of the experimental framework, providing details on specific implementations, evaluation benchmarks, and MLLMs used for comparative evaluation. Following this, we conduct a comprehensive comparison of Parrot with the state-of-the-art approaches using multilingual benchmarks. We also compare Parrot with leading models across a range of multimodal tasks. Finally, we conclude with ablation studies and visualization of multilingual cases, highlighting the exceptional ability of Parrot in handling multilingual tasks.
Table 1: Accuracy performance comparison on multilingual benchmarks. We report all compared methods with VLMEvalKit (Duan et al., 2024). The best and second results are shown in bold and underline, respectively.
@@
VisualGLM (Du et al., 2022) ChatGLM-6B 31.05 18.07 19.42 15.38 22.81 19.77 23.2 17.18 11.43 2.92 6.62 5.33
VisCPM-Chat (Hu et al., 2023) CPM-Bee-10B 53.10 47.54 28.19 26.90 26.78 26.84 45.88 46.39 15.81 1.46 9.19 1.20
Qwen-VL-Chat (Bai et al., 2023b) Qwen-7B 56.02 57.77 46.37 43.04 41.05 48.65 54.29 56.52 43.12 35.73 39.17 42.86
InstructBLIP (Dai et al., 2023) Vicuna-7B 39.47 32.92 35.67 23.80 28.36 36.37 27.83 18.81 27.14 3.26 8.50 20.87
mPLUG-Owl2 (Ye et al., 2023) LLaMA2-7B 67.25 60.99 59.70 45.78 45.43 62.63 66.15 59.36 58.24 37.88 47.68 60.39
Monkey (Li et al., 2023d) Qwen-VL-7B 66.02 58.18 46.31 38.83 37.66 48.59 58.07 53.52 49.57 31.01 31.35 45.18
LLaVA-1.5 (Liu et al., 2023a) Vicuna-v1.5-7B 67.07 58.83 59.76 43.50 46.43 59.06 65.37 58.33 59.02 36.16 43.90 56.95
LLaVA-NeXT (Liu et al., 2024) LLaMA3-8B 70.92 64.33 63.21 48.34 48.02 66.35 69.80 63.31 61.83 47.64 47.03 64.99
DeepSeek-VL-7B (Lu et al., 2024a) Deepseek-7B 72.66 65.95 64.41 49.70 49.07 67.57 70.71 64.03 62.61 48.05 47.95 65.53
GLM-4v-9B (GLM et al., 2024) GLM-4-9B 69.26 62.83 61.59 47.23 46.9 64.35 67.91 61.31 60.01 46.13 45.7 63.09
ShareGPT4V (Chen et al., 2023b) Vicuna-v1.5-7B 69.24 60.23 60.29 43.57 45.26 61.23 69.59 61.6 59.62 37.37 43.38 59.45
Idefics3-8B (Laurençon et al., 2024) LLaMA3-8B 74.50 67.70 66.08 50.91 50.41 69.66 73.50 66.79 65.53 49.85 49.79 68.68
Qwen2-VL (Wang et al., 2024) Qwen2-7B 80.51 80.23 78.11 74.07 71.72 79.33 78.59 78.39 75.93 74.70 73.45 75.88
LLaVA-OneVision (Li et al., 2024b) Qwen2-7B 79.03 78.23 75.91 73.36 67.79 76.37 76.74 75.30 73.45 70.44 64.85 73.14
@@
Parrot Qwen2-7B 80.11 80.03 79.62 76.55 75.02 79.94 78.02 77.16 76.76 75.92 74.05 77.71
4.1 Experimental Setup
Implementation Details. In this study, we configure Parrot with the pre-trained CLIP ViT-L/14 (Radford et al., 2021) as the vision encoder. To validate the effectiveness of Parrot, we select both Qwen1.5 and Qwen2 (Bai et al., 2023a) as the backbones. The initial learning rates for the two stages are set at $1e^{-3}$ and $2e^{-5}$, respectively, with the batch size of 256 and 128. The entire training process is optimized to 21 hours on the 16$\times$×A100 GPUs setup, benefiting from the relatively small training datasets. Additionally, BF16 and TF32 precision formats are employed to balance speed and accuracy throughout the training process. As defined in Eq. 4, we set the number of experts to six to correspond with the number of languages. Each expert is an MLP composed of two linear layers with SiLU (Elfwing et al., 2018) activation function. More details are provided in Table 13.
Evaluation Benchmark. Our evaluation consists of two parts: one assessing the multilingual capabilities of MLLMs, while the other evaluating its overall performance. The first part is conducted on two datasets: multilingual MMBench (Liu et al., 2023c) and a newly developed benchmark MMMB. The second part of the evaluation spans a wide range of multimodal tasks, such as MME (Fu et al., 2023), MMStar (Chen et al., 2024), ScienceQA (Lu et al., 2022), RealWorldQA (x.ai, 2024) and SEED-Bench (Li et al., 2024a), with performance visualized in a radar chart in Figure 7.
Figure 6: t-SNE visualizations of LLaVA and Parrot using different vision encoders.
Comparison Models. For comprehensive comparisons, we select leading open-source models in MLLMs, including LLaVA-1.5 (Li et al., 2023a), LLaVA-NeXT (Liu et al., 2024), Monkey (Li et al., 2023d), VisualGLM (Du et al., 2022), VisCPM (Hu et al., 2023), GLM-4v (GLM et al., 2024), ShareGPT4V (Chen et al., 2023b), InstructBLIP (Dai et al., 2023), mPLUG-Owl2 (Ye et al., 2023), Idefics3 (Laurençon et al., 2024), LLaVA-OneVision (Li et al., 2024b), and Qwen2-VL (Wang et al., 2024). For the evaluation process, we employ the VLMEvalKit (Duan et al., 2024) in OpenCompass, ensuring consistent configuration settings across all methods to maintain fairness in comparison.